# Invariant subspace perturbations related to defective eigenvalues of $\Delta$-Hermitian and Hamiltonian matrices

### v1

@@ 2 Preliminaries
The following result shows the perturbation behaviors of the invariant subspaces corresponding to part of the eigenvalues of $\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)$. Suppose $F\in{\mathbb{C}}^{n\times\ell}$ is a full column rank matrix and
@@
$F=\left[\begin{array}[]{cccc}Q^{T}&(Q\Omega)^{T}&\ldots&(Q\Omega^{\rho-1})^{T}\end{array}\right]$ (6)
@@
$X_{\Omega}(t)=\left[\begin{array}[]{ccccc}X_{\Omega,1}^{T}(t)&\ldots&X_{\Omega,\rho}^{T}(t)&\ldots&X_{\Omega,m}^{T}(t)\end{array}\right]+O(t),$
$X_{\Omega,k}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}&k<\rho\\ \left[\begin{array}[]{cccc}Q^{T}&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad\quad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ \left[\begin{array}[]{cccc}(G_{k\rho}Q)^{T}&0&\ldots&0\end{array}\right]^{T}&\\ \qquad+O(\left[\begin{array}[]{cccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t^{1-1/\rho}&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
Proof. See, e.g., [12, 20, 27].
@@
Proof. See, e.g., [10, 15, 18].
@@ 3 Invariant subspace perturbations of $\Delta$-Hermitian matrices
Similarly, there is $Z_{2}(t)=I+O(t^{2})$ such that
@@
$(C+tD)UZ(t)=UZ(t)\left[\begin{array}[]{cc}\widetilde{C}_{T}(t)&0\\ 0&\widetilde{C}_{C}(t)\end{array}\right]$
@@
$(UZ(t))^{\ast}\Delta(UZ(t))=Z^{\ast}(t)\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]Z(t)=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right].$
@@ 3.1 Nonreal eigenvalue case
Since $P^{\ast}M_{11}^{\ast}P=P^{\ast}V^{\ast}(\Delta D)V_{c}P$, the matrix corresponds to $W_{k}$ in (2) is
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
@@
$\widetilde{V}_{\rho}(t)=UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]\\ 0\end{array}\right],$ (52)
@@
$(C+tD)\widehat{V}_{\rho}^{(c)}(t)=\widehat{V}_{\rho}^{(c)}(t)(P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho})$
@@
$X^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{s_{\rho}}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}}),&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@
$\widehat{X}^{(c)}_{k\rho}(t)=X^{(c)}_{k\rho}(t)R_{\rho}^{-1}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{\rho}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})&\\ \qquad+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
@@
The next result is about the perturbations of the invariant subspaces corresponding to part of the eigenvalues of $\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t)$ and $(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast}.$
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho\end{array}\right$.
@@
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@
Under the conditions of Theorem 3, suppose $\omega$ is a simple eigenvalue of $\Theta_{\rho}(S_{\rho})$.
@@
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T}),&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T}),&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T}),&k>\rho.\end{array}\right$.
@@
In Corollary 5, $l_{\rho}(t)$ and $l_{\rho}^{(c)}(t)$ are right and left eigenvectors of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{w}(t)$. $l_{\rho}(t),l_{\rho}^{(c)}(t)=O(1)$ and $l_{\rho}^{\ast}(t)l_{\rho}^{(c)}(t)=t^{1-1/\rho}$. The latter indicates that the spectral condition numbers of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{\omega}(t)$ is $O(t^{-(1-1/\rho)})$. If we consider $C$ is the matrix perturbed from $C+tD$, since the spectral condition number is $O(t^{-(1-1/\rho)})$, the eigenvalue should change with an order $O(t^{1/\rho})$. In this sense, Corollary 5 generalizes the standard first order perturbation results for simple eigenvalues ([13, Sec.7.2.2]).

### v2

@@ 2 Preliminaries
The following result shows the perturbation behaviors of the invariant subspaces corresponding to a subset of $\Lambda(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))$. Suppose $F\in{\mathbb{C}}^{n\times\ell}$ is a full column rank matrix and
@@
$X_{\Omega}(t)=\left[\begin{array}[]{ccccc}X_{\Omega,1}^{T}(t)&\ldots&X_{\Omega,\rho}^{T}(t)&\ldots&X_{\Omega,m}^{T}(t)\end{array}\right]^{T}+O(t),$
$X_{\Omega,k}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}Q^{T}&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}&\\ \qquad\qquad\qquad\qquad\qquad\quad+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ \left[\begin{array}[]{cccc}(G_{k\rho}Q)^{T}&0&\ldots&0\end{array}\right]^{T}&\\ \qquad+O(\left[\begin{array}[]{cccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t^{1-1/\rho}&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
Proof. This is a slightly modified form of those in [12, 20, 27].
@@
Proof. This is a slightly modified form of those in [10, 15, 18].
@@ 3 Invariant subspace perturbations of $\Delta$-Hermitian matrices
Similarly, there is a $Z_{2}(t)=I+O(t^{2})$ such that
@@
$(C+tD)UZ(t)=UZ(t)\left[\begin{array}[]{cc}\widetilde{C}_{T}(t)&0\\ 0&\widetilde{C}_{C}(t)\end{array}\right]$ (18)
@@
$(UZ(t))^{\ast}\Delta(UZ(t))=Z^{\ast}(t)\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right]Z(t)=\left[\begin{array}[]{cc}\Delta_{T}&0\\ 0&\Delta_{C}\end{array}\right].$ (24)
@@ 3.1 Nonreal eigenvalue case
Since $P^{\ast}M_{11}^{\ast}P=P^{\ast}V^{\ast}(\Delta D)V_{c}P$, the matrix corresponding to $W_{k}$ in (2) is
@@
    $\widetilde{X}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(0_{s_{\rho}},t^{1/\rho}I_{s_{\rho}},\ldots,t^{1-1/\rho}I_{s_{\rho}})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
    $\widetilde{X}^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
$\widetilde{V}_{\rho}(t)=UZ(t)\left[\begin{array}[]{c}\widetilde{Z}_{1}(t)\left[\begin{array}[]{c}X_{\rho}(t)\\ 0\end{array}\right]\\ 0\end{array}\right]$ (52)
@@
$(C+tD)\widehat{V}_{\rho}^{(c)}(t)=\widehat{V}_{\rho}^{(c)}(t)(P_{\rho}^{\ast}(\bar{\lambda}I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}^{(c)}(t))P_{\rho}),$
@@
$X^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{s_{\rho}}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@
$\widehat{X}^{(c)}_{k\rho}(t)=X^{(c)}_{k\rho}(t)R_{\rho}^{-1}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&I_{\rho}\end{array}\right]+\mathop{\rm diag}\nolimits(t^{1-1/\rho}I_{s_{\rho}},\ldots,t^{1/\rho}I_{s_{\rho}},0_{s_{\rho}})&\\ \qquad+O(\left[\begin{array}[]{cccc}t&t^{1-1/\rho}&\ldots&t^{1/\rho}\end{array}\right]^{T}),&k=\rho\\ \left[\begin{array}[]{cc}0&0\\ 0&G_{k\rho}^{(c)}\end{array}\right]+O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho,\end{array}\right$.
@@
The next result is about the perturbations of the invariant subspaces corresponding to subsets of $\Lambda(\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))$ and $\Lambda((\lambda I_{\rho s_{\rho}}+t^{1/\rho}\widetilde{\Theta}_{\rho}(t))^{\ast})$.
@@
$L_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&(Q(t^{1/\rho}\Omega))^{T}&\ldots&(Q(t^{1/\rho}\Omega)^{\rho-1})^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho\end{array}\right$.
@@
$L^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\Omega)^{\rho-1}(Q^{(c)})^{\ast}&\ldots&(t^{1/\rho}\Omega)(Q^{(c)})^{\ast}&0\end{array}\right]^{\ast}&\\ \qquad\qquad\qquad\qquad\qquad\qquad+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@
Under the conditions of Theorem 5, suppose $\omega$ is a simple eigenvalue of $\Theta_{\rho}(S_{\rho})$. The vectors
@@
$l_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-k/\rho}&t^{1-(k-1)/\rho}&\ldots&t^{1-1/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}0&t^{1/\rho}\omega q^{T}&\ldots&(t^{1/\rho}\omega)^{\rho-1}q^{T}\end{array}\right]^{T}+O(\left[\begin{array}[]{cccc}t^{1/\rho}&t^{2/\rho}&\ldots&t\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t^{1/\rho}&t^{1/\rho}&t^{2/\rho}&\ldots&t&\ldots&t\end{array}\right]^{T})&k>\rho,\end{array}\right$.
$l^{(c)}_{k\rho}(t)=\left\{\begin{array}[]{ll}O(\left[\begin{array}[]{cccc}t^{1-1/\rho}&t^{1-2/\rho}&\ldots&t^{1-k/\rho}\end{array}\right]^{T})&k<\rho\\ \left[\begin{array}[]{cccc}(t^{1/\rho}\omega)^{\rho-1}(q^{(c)})^{\ast}&\ldots&t^{1/\rho}\omega(q^{(c)})^{\ast}&0\end{array}\right]^{\ast}+O(\left[\begin{array}[]{cccc}t&\ldots&t^{2/\rho}&t^{1/\rho}\end{array}\right]^{T})&k=\rho\\ O(\left[\begin{array}[]{ccccccc}t&\ldots&t&\ldots&t^{2/\rho}&t^{1/\rho}&t^{1/\rho}\end{array}\right]^{T})&k>\rho.\end{array}\right$.
@@
In Corollary 7, $l_{\rho}(t)$ and $l_{\rho}^{(c)}(t)$ are right and left eigenvectors of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{w}(t)$. Note that $l_{\rho}(t),l_{\rho}^{(c)}(t)=O(1)$ and $l_{\rho}^{\ast}(t)l_{\rho}^{(c)}(t)=t^{1-1/\rho}$. The latter indicates that the spectral condition numbers of $C+tD$ corresponding to the eigenvalue $\lambda+t^{1/\rho}\widetilde{\omega}(t)$ is $O(t^{-(1-1/\rho)})$. If we consider $C$ as the matrix perturbed from $C+tD$, since the spectral condition number is $O(t^{-(1-1/\rho)})$, the eigenvalue should change with an order $O(t^{1/\rho})$. In this sense, Corollary 7 generalizes the standard first order perturbation results for simple eigenvalues ([13, Sec.7.2.2]).